OpenStudy (anonymous):

Given that sin A = 1/3 and angle A is in the first quadrant, find cos A.

OpenStudy (anonymous):

\[\cos A=\sqrt{1-\sin ^2A}\]

OpenStudy (anonymous):

sooo (1/3)^2 -1

OpenStudy (anonymous):

http://screencast.com/t/YM8G3sGx

OpenStudy (anonymous):

1 - (1/3)

OpenStudy (anonymous):

\[\sqrt{1-\left( \frac{ 1 }{ 3 } \right)^2}=\sqrt{1-\frac{ 1 }{ 9 }}=\sqrt{\frac{ 8 }{ 9 }}=\frac{ 2\sqrt{2} }{ 3 }\]

OpenStudy (anonymous):

ah yes

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

also.. is tan (3π/2) =

OpenStudy (anonymous):

is the solution undefined?